# Quantum combinatorial optimization beyond the variational paradigm: simple schedules for hard problems

### v1

@@ I Introduction
Our main contribution to the questions we have raised is the following. By combining the geometric approach to adiabatic quantum computation (AQC) [19] with the spin coherent-state path integral for quantum spin glasses [20, 21, 22], we show in detail how semi-classical observables can be used to systematically derive “adaptive” annealing schedules at $\mathrm{poly}(N)$ overhead, where $N$ is the problem size. Such schedules can lead to orders-of-magnitude improvements in the AQC probability $P$ to observe the optimal solution of a given problem instance. We work with two established hard-instance data sets: one for the Sherrington-Kirkpatrick (SK) model as used in Ref. [22], the other for Maximum 2-Satisfiability (MAX 2-SAT) as introduced earlier by Shor et al. [15]. On the latter data, we improve the geometric mean success probability across all instances from $0.01\%$ in the linear-schedule case to $3.1\%$ with our method; for single instances, the improvement can reach more than four orders of magnitude (cf. section V.3). For the SK model, we find that our schedules are close to the “ideal” ones derived from the exact adiabatic bottleneck.
@@ II Quantum Combinatorial Optimization
For the purposes of this work, we do not differentiate between AQC and the QAOA in the sense that one may consider fixed-angle QAOA the Trotterized version of continuous-schedule AQC. Therefore, after defining the continuous-time adiabatic Hamiltonian below, we show once how to initialize the corresponding QAOA angles based on a given annealing protocol. In our theoretical derivations, we then stick to a formulation in terms of continuous time, while we effectively use fixed-angle QAOA in our numerics.
@@ II.1 Problem Definition
$\beta_{k}=\tau s^{1}(k/p),\quad\gamma_{k}=\tau s^{2}(k/p),\quad\tau=T/p.$ (5)
@@ II.3 Gaussian Fluctuations
to define the corresponding quasi-particles [22] with creation and annihilation operators $\hat{\eta}_{i}^{{\dagger}}$, $\hat{\eta}_{i}^{\phantom{\dagger}}$. Since Hamiltonians in the context of combinatorial optimization typically have frustrated ground states, these quasi-particles are usually called paramagnons.
@@
Together with the initial condition $\bm{F}(t,t)=-\mathrm{i}\sigma_{3}$, Eq. (17) defines a $2N\times 2N$ initial-value problem. Regarding the algorithmic overhead of our method, this $\sim N^{2}$ cost is the biggest contribution since the number of mean-field equations of motion (7) scales as $\sim N$. Once Eq. (17) is solved, the additional overhead of obtaining the corresponding annealing schedules, which we describe in detail in section III, is actually independent of the system size $N$. In total, then, we have an algorithmic overhead that scales as $\mathrm{poly}(N)$.
@@ III Geometric Approach to Optimal Schedules
where the prime denotes differentiation. Note that due to these, Eq. (23) is not an initial-value problem but rather a second-order boundary-value problem (BVP), which can be solved numerically, e.g., by means of the monotonic implicit Runge-Kutta (MIRK) or the shooting method [28, 29].
For hard instances, the one-parameter metric “tensor” will almost always be dominated strongly by the energy gap, i.e. we can assume
@@ III.1 Adaptive Schedules
which can be thought of as indicating the degree of localization of a given classical trajectory. An example of the semi-classical paramagnon numbers $\mathcal{N}_{i}(s)$ is shown in the upper panel of Fig. 3. As was already shown in Ref. [22], the paramagnons of spins with high frustration (cf. Fig. 1) tend to grow around the bottleneck $s_{*}$, which one can understand physically by associating it with the finite-size analog of a first-order phase transition [3]. If a frustrated spin tries to localize and starts to fluctuate strongly while doing so, this turns out to be a very good semi-classical indicator of the adiabatic bottleneck, which we quantify by combining the local Edwards-Anderson order parameter with the paramagnon number to form the so-called localization susceptibility
@@
follows from the stereographic projection of the Bloch sphere to the complex plane [21, 20]. We remark that the behavior seen in the lower panel of Fig. 3 is indeed typical [22], thus providing us with a method to replace the exact bottleneck (26) with an approximation $\tilde{s}_{*}$ derived from the maxima of $\chi_{i}(s)$. Note that multiple peaks as visible in Fig. 3 do not pose fundamental a problem to this strategy, as one may simply perform several quantum circuits for each bottleneck “candidate”; in most cases, also taking $\mathcal{N}_{i}(s)$ into direct consideration will render this unnecessary.
@@ IV.1 Sherrington-Kirkpatrick
To extend our scaling results to larger system sizes, to these data we add 20 hard instances at $N=22$ that have been found by the following procedure: we run the mean-field evolution for a large number of random instances ($\sim 20000$) and retain only those for which the final effective magnetization of at least one spin is close to zero (the reasoning behind this is that we expect frustration and hardness to be correlated); then we run linear AQC for $T=2^{7}$ and keep those instances for which the success probability is below half a percentage (similar to the strategy Shor et al. [15]). As is indicated by our results in section V, this procedure seems to produce even harder instances than those contained in the remainder of the SK data set.
@@
Table 1: Ensemble sizes for the hard SK data set. Note that the 20 instances at $N=22$ are found by a different procedure than the smaller instances (s. main text).
@@ V Results
Digitizing such analog quantum evolution by Trotterization into many QAOA layers via Eq. (5), which then have to be run on a digital device, is likely still challenging even at present-day improved fidelities. For this reason, in units where $h_{X}=1$, the maximal final time we consider is $T=1024$. To obtain an order-of-magnitude intuition for the success probabilities we encounter below, consider a typical low linear-schedule success probability of $P_{\mathrm{lin}}=10^{-6}$. Then even $R=10^{4}$ runs of the quantum device will result in a failure probability (i.e. the probability of not seeing the optimal solution) of about $(1-P_{\mathrm{lin}})^{R}\approx 99\%$. Since $P_{\mathrm{lin}}$ is bound to become worse for larger systems sizes $N$, one can already see from this simple back-of-the-envelope estimate why it is difficult to optimize annealing schedules just by trial and error: if the ground state is “invisible” for some finite $R$, then it will effectively not influence the optimization.
Before coming to the presentation of our main results in sections V.2 and V.3, in the next section we first discuss an observation that was already made by Shor et al. [15], namely that shorter evolution times $T$, in fact, seem to increase the success probability $P$, which they referred to as “The hare beats the tortoise”. The obvious question to ask is whether this is simply a finite-size effect at small $N$.
V.1 Is diabatic enhancement a finite-size effect?
@@
To conclude this section, we turn to Fig. 6, which shows the ensemble statistics of the most likely final states $|\lambda\rangle$, computed for $T=128$ with the full SK ensembles as quoted in Tab. 1. For one thing, the fraction of instances for which the ground state is most likely turns out to be very small (which is, of course, to be expected by construction). More strikingly, there is a clear trend with $N$ towards higher excited states. We believe this should be interpreted as a clear indicator that diabatic enhancement is unlikely to be useful for large systems. To put it concisely, if it were otherwise and simply running fast anneals for arbitrary problems presented a generically good strategy, it would appear that we had found a means to solve NP-hard problems efficiently.
V.2 Sherrington-Kirkpatrick
Having come to the conclusion that fast anneals most likely do not present a sustainable strategy, we move on to discuss our main results for the SK data set, shown in Fig. 7 and 8. The effective ensemble sizes after imposing a success-probability cutoff $P<0.5\%$ are given in Tab. 2. We use this cutoff to make sure our analysis is focused on instances that are actually hard for the linear schedules at each $T$. Since it is numerically very expensive to find such instances, the ensemble sizes are moderate.
@@ V.2 Sherrington-Kirkpatrick
The “ideal” results shown in both Fig. 7 and 8 are calculated by inserting the exact bottleneck position $s_{*}$ into Eq. (32), whereas the approximate $\tilde{s}_{*}$ is used for the “adaptive” results; for example schedules, s. Fig. 4. Note also that we only know the exact bottleneck $s_{*}$ to a relatively low precision of $2^{-5}$, i.e. one should not expect our “ideal” results to be fully ideal; for $N=22$, we did not attempt to compute the exact $s_{*}$ and hence do not have the corresponding “ideal” data point.
One of the most striking features of Fig. 7 is that the geometric mean $\bar{P}$ across the ensembles goes down systematically with $T$ when using a linear schedule. In light of our previous discussion (section V.1) of diabatic enhancement, we do not conclude from this that running for shorter times is a valid strategy. A related obvious question to ask is whether there is anything to learn from the fact that the decrease in $\bar{P}$ is stronger for larger $N$? One hypothesis consistent with the data appears to be this: the smaller instances (i.e. $N=8,...,14$) are, in fact, evolving almost adiabatically already for $T=128$, while the larger (and statistically harder) instances cross over into the adiabatic regime from left to right. Given Fig. 6, we argue that the results we see for $N>14$ in the two right panels should inform our generic expectations: in an energy landscape with many local minima far away (in Hamming distance) from the optimum, a return to the ground state after excitation cannot be relied upon.
@@
V.3 Maximum 2-Satisfiability
While the results on the SK instances are very promising, one might argue that the SK model is bound to benefit from mean-field models by its very nature. To guard against this possibility, we now move on to the investigation of the hard MAX 2-SAT instances from Shor et al. [15], which have very uniform local fields $h_{i}$ and interactions $J_{ij}$. An exemplary set of mean-field trajectories was already shown above in Fig. 5. Note that it is typical for MAX 2-SAT to possess a greater number of non-localizing spins, i.e. spins that have $n_{i}^{z}(1)\approx 0.0$. Interestingly, while this means that the mean-field approximate optimization algorithm [21] results in a toin coss for these spins in particular, their semi-classical fluctuations contain useful information about the adiabatic bottleneck (cf. lowest panel of Fig. 5).
@@ V.3 Maximum 2-Satisfiability
The main results of this section, obtained from running AQC (or, for that matter, fixed-angle QAOA) simulations for all instances with the three different types of schedules, are given in Fig. 10 and 11. Note again that we employ a second-order Trotter-Suzuki expansion [24] for the linear schedules. We have also verified for $T=128$ that lowering the step size from $T/p=2^{-3}$ to $2^{-4}$ does indeed slightly improve our adaptive results, i.e. the results obtained at the former step size for $T=256$ should be conservative.
Figure 11: Histograms detailing the same results as shown in Fig. 10. Observe the reduce variance in the “effective” results, an explanation for which is provided in the main text. For $T=128$ (upper panel), the geometric mean success probabilities are $\bar{P}_{\mathrm{lin}}=0.007\%$, $\bar{P}_{\mathrm{ad}}=1.1\%$ and $\bar{P}_{\mathrm{eff}}=1.2\%$ for the linear, adaptive and effective schedules, respectively; for $T=256$ (lower panel), we instead find $\bar{P}_{\mathrm{lin}}=0.01\%$, $\bar{P}_{\mathrm{ad}}=2.3\%$ and $\bar{P}_{\mathrm{eff}}=3.1\%$.
@@
While the linear results hardly improve from top to bottom in Fig. 11, the adaptive results visibly shift towards higher success probabilities. Since Shor et al. [15] also did not observe an increase in the linear success probability for larger final times than $T=100$, it seems possible to interpret the fact that our results do improve from top to bottom as an indicator that we indeed predict the true bottleneck reasonably accurately. The lower spread in the “effective” results, stemming from the corresponding schedule shown in Fig. 9, is also worth commenting on: apparently, the instances ending up in the left tails of the “adaptive” distributions have indeed received below-average $\tilde{s}_{*}$ predictions that are then corrected by using $\tilde{s}_{*}=0.62$ instead; however, it would appear that the opposite happens on the right tail where presumably very accurate predictions are shifted towards lower probabilities.

### v2

@@ I Introduction
Practical applications of discrete optimization are manifold and have been widely discussed in the literature [19, 20]. Prominent examples are use cases in financial [21] and energy markets [22, 23]. Hard problems in biology such as protein folding [24] have also been addressed.
Our main contribution to the questions we have raised is the following. By combining the geometric approach to adiabatic quantum computation (AQC) [25] with the spin coherent-state path integral for quantum spin glasses [26, 27, 28], we show in detail how semi-classical observables can be used to systematically derive “adaptive” annealing schedules at $\mathrm{poly}(N)$ overhead, where $N$ is the problem size. Such schedules can lead to orders-of-magnitude improvements in the AQC probability $P$ to observe the optimal solution of a given problem instance. We work with two established hard-instance data sets with known exact ground states: one for the Sherrington-Kirkpatrick (SK) model as used in Ref. [28], the other for Maximum 2-Satisfiability (MAX 2-SAT) as introduced earlier by Shor et al. [15]. On the latter data, we improve the geometric mean success probability across all instances from $0.01\%$ in the linear-schedule case to $3.1\%$ with our method; for single instances, the improvement can reach more than four orders of magnitude (cf. section V.2). For the SK model, we find that our schedules are close to the “ideal” ones derived from the exact adiabatic bottleneck.
@@ II Quantum Combinatorial Optimization
For the purposes of this work, we do not differentiate between AQC and the QAOA in the sense that one may consider the Trotterized version of continuous-schedule AQC a limiting case of fixed-angle QAOA. Therefore, after defining the continuous-time adiabatic Hamiltonian below, we show once how to initialize the corresponding QAOA angles based on a given annealing protocol. In our theoretical derivations, we then stick to a formulation in terms of continuous time, while we effectively use fixed-angle QAOA in our numerics.
@@ II.1 Problem Definition
$\beta_{k}=\tau s^{1}(Tk/p),\quad\gamma_{k}=\tau s^{2}(Tk/p),\quad\tau=T/p.$ (5)
@@ II.3 Gaussian Fluctuations
to define the corresponding quasi-particles [28] with creation and annihilation operators $\hat{\eta}_{i}^{{\dagger}}$, $\hat{\eta}_{i}^{\phantom{\dagger}}$. Since Hamiltonians in the context of combinatorial optimization typically have frustrated ground states, their quasi-particles are usually called paramagnons (as opposed to, e.g., the quasi-particle excitations on top of a ferromagnetically ordered ground state, which would be addressed simply as magnons).
@@
Together with the initial condition $\bm{F}(0,0)=-\mathrm{i}\sigma_{3}$, Eq. (17) defines a $2N\times 2N$ initial-value problem. Regarding the algorithmic overhead of our method, this $\sim N^{2}$ cost is the biggest contribution since the number of mean-field equations of motion (7) scales as $\sim N$. Once Eq. (17) is solved, the additional overhead of obtaining the corresponding annealing schedules, which we describe in detail in section III, is actually independent of the system size $N$. In total, then, we have an algorithmic overhead that scales as $\mathrm{poly}(N)$.
@@ III Geometric Approach to Optimal Schedules
where $g(s)$ is a one-parameter metric “tensor” and the prime denotes differentiation. Note that due to these, Eq. (23) is not an initial-value problem but rather a second-order boundary-value problem (BVP), which can be solved numerically, e.g., by means of the monotonic implicit Runge-Kutta (MIRK) or the shooting method [34, 35].
For hard instances, the metric will almost always be dominated strongly by the energy gap, i.e. we can assume
@@ III.1 Adaptive Schedules
which can be thought of as indicating the degree of localization of a given classical trajectory. An example of the semi-classical paramagnon numbers $\mathcal{N}_{i}(s)$ is shown in the upper panel of Fig. 3. As was already shown in Ref. [28], the paramagnon number of spins with high frustration (cf. Fig. 1) tends to grow around the bottleneck $s_{*}$, which one can understand physically by associating it with the finite-size analog of a first-order phase transition [3]. If a frustrated spin tries to localize and starts to fluctuate strongly while doing so, this turns out to be a very good semi-classical indicator of the adiabatic bottleneck, which we quantify by combining the local Edwards-Anderson order parameter with the paramagnon number to form the so-called localization susceptibility
@@
follows from the stereographic projection of the Bloch sphere to the complex plane [27, 26]. We remark that the behavior seen in the lower panel of Fig. 3 is indeed typical [28], thus providing us with a method to replace the exact bottleneck (26) with an approximation $\tilde{s}_{*}$ derived from the maxima of $\chi_{i}(s)$. Note that multiple peaks as visible in Fig. 3 do not pose fundamental a problem to this strategy, as one may simply perform several quantum circuits for each bottleneck “candidate”; in most cases, also taking $\mathcal{N}_{i}(s)$ into direct consideration will render this unnecessary, in the sense that the degree of frustration of a particular spin trajectory contains information about its relevance for the bottleneck.
@@ IV.1 Sherrington-Kirkpatrick
To extend our scaling results to larger system sizes, to these data we add 20 hard instances at $N=22$ that have been found by the following procedure: we run the mean-field evolution for a large number of random instances ($\sim 20000$) and retain only those for which the final effective magnetization of at least one spin is close to zero (the reasoning behind this is that we expect frustration and hardness to be correlated); then we run linear AQC for $T=2^{7}$ and keep those instances for which the success probability is below half a percentage (similar to the strategy Shor et al. [15]). As is indicated by our results in section V, this procedure seems to produce even harder instances than those contained in our previous SK data set [28], where the selection of hard instances was based entirely on the size of the spectral gap obtained from exact diagonalization of the quantum Hamiltonian.
@@
Table 1: Ensemble sizes for the hard SK data set. As the computational cost for the largest system size $N=22$ is already considerable, we are limited to a relatively small number of instances.
@@ V Results
Digitizing such analog quantum evolution by Trotterization into many QAOA layers via Eq. (5), which then have to be run on a digital device, is likely still challenging even at present-day improved fidelities. For this reason, in units where $h_{X}=1$, the maximal final time we consider is $T=1024$. To obtain an order-of-magnitude intuition for the success probabilities we encounter below, consider a typical low linear-schedule success probability of $P_{\mathrm{lin}}=10^{-6}$. Then even $R=10^{4}$ runs of the quantum device will result in a failure probability (i.e. the probability of not seeing the optimal solution) of about $(1-P_{\mathrm{lin}})^{R}\approx 99\%$. Since $P_{\mathrm{lin}}$ is bound to become worse for larger systems sizes $N$, one can already see from this simple back-of-the-envelope estimate why it is difficult to optimize annealing schedules for maximal success probability just by trial and error: if the ground state is “invisible” for some finite $R$, then it will effectively not influence the optimization. Note, however, that a choice of cost function other than the success probability might not suffer from this limitation.
@@
V.1 Sherrington-Kirkpatrick
Our main results for the SK data set are shown in Fig. 6 and 7. The effective ensemble sizes after imposing a success-probability cutoff $P<0.5\%$ are given in Tab. 2. We use this cutoff to make sure our analysis is focused on instances that are actually hard for the linear schedules at each $T$. Since it is numerically very expensive to find such instances, the ensemble sizes are moderate.
@@ V.1 Sherrington-Kirkpatrick
The “ideal” results shown in both Fig. 6 and 7 are calculated by inserting the exact bottleneck position $s_{*}$ into Eq. (32), whereas the approximate $\tilde{s}_{*}$ is used for the “adaptive” results; for example schedules, s. Fig. 4. Note also that we only know the exact bottleneck $s_{*}$ to a relatively low precision of $2^{-5}$, i.e. one should not expect our “ideal” results to be fully ideal; for $N=22$, we did not attempt to compute the exact $s_{*}$ and hence do not have the corresponding “ideal” data point.
One of the most striking features of Fig. 6 is that the geometric mean $\bar{P}$ across the ensembles goes down systematically with $T$ when using a linear schedule. As discussed in detail in Appendix .1, we do not conclude from this that running for shorter times is a valid strategy. A related obvious question to ask is whether there is anything to learn from the fact that the decrease in $\bar{P}$ is stronger for larger $N$? One hypothesis consistent with the data appears to be this: the smaller instances (i.e. $N=8,...,14$) are, in fact, evolving almost adiabatically already for $T=128$, while the larger (and statistically harder) instances cross over into the adiabatic regime from left to right. Given Fig. 11, we argue that the results we see for $N>14$ in the two right panels should inform our generic expectations: in an energy landscape with many local minima far away (in Hamming distance) from the optimum, a return to the ground state after excitation cannot be relied upon.
@@
V.2 Maximum 2-Satisfiability
@@
While the results on the SK instances are very promising, one might argue that the SK model is bound to benefit from mean-field models by its very nature. To guard against this possibility, we now move on to the investigation of the hard MAX 2-SAT instances from Shor et al. [15], which have very uniform local fields $h_{i}$ and interactions $J_{ij}$. An exemplary set of mean-field trajectories was already shown above in Fig. 5. Note that it is typical for MAX 2-SAT to possess a greater number of non-localizing spins, i.e. spins that have $n_{i}^{z}(1)\approx 0.0$. Interestingly, while this means that the mean-field approximate optimization algorithm [27] results in a coin toss for these spins in particular, their semi-classical fluctuations contain useful information about the adiabatic bottleneck (cf. lowest panel of Fig. 5).
@@
The main results of this section, obtained from running AQC (or, for that matter, fixed-angle QAOA) simulations for all instances with the three different types of schedules, are given in Fig. 9 and 10. Note again that we employ a second-order Trotter-Suzuki expansion [30] for the linear schedules. We have also verified for $T=128$ that lowering the step size from $T/p=2^{-3}$ to $2^{-4}$ does indeed slightly improve our adaptive results, i.e. the results obtained at the former step size for $T=256$ should be conservative.
@@
While the linear results hardly improve from top to bottom in Fig. 10, the adaptive results visibly shift towards higher success probabilities. Since Shor et al. [15] also did not observe an increase in the linear success probability for larger final times than $T=100$, it seems possible to interpret the fact that our results do improve from top to bottom as an indicator that we indeed predict the true bottleneck reasonably accurately. The lower spread in the “effective” results, stemming from the corresponding schedule shown in Fig. 8, is also worth commenting on: apparently, the instances ending up in the left tails of the “adaptive” distributions have indeed received below-average $\tilde{s}_{*}$ predictions that are then improved by using the aggregate value $\tilde{s}_{*}=0.62$ instead; however, it would appear that the opposite happens on the right tail where presumably very accurate predictions are shifted towards lower probabilities.
@@ VI Conclusion
Figure 10: Histograms detailing the same results as shown in Fig. 9. Observe the reduced variance in the “effective” results, an explanation for which is provided in the main text. For $T=128$ (upper panel), the geometric mean success probabilities are $\bar{P}_{\mathrm{lin}}=0.007\%$, $\bar{P}_{\mathrm{ad}}=1.1\%$ and $\bar{P}_{\mathrm{eff}}=1.2\%$ for the linear, adaptive and effective schedules, respectively; for $T=256$ (lower panel), we instead find $\bar{P}_{\mathrm{lin}}=0.01\%$, $\bar{P}_{\mathrm{ad}}=2.3\%$ and $\bar{P}_{\mathrm{eff}}=3.1\%$.
@@
Note also that, while solving the embedding problem, i.e. the problem of mapping a given discrete optimization problem onto an existing quantum hardware graph, is a challenge in itself [41, 42], our tools are embedding-agnostic and can be leveraged to improve the performance arbitrary hardware configurations.
Further interesting avenues could be found in applying our methods to improve alternative combinatorial-optimization frameworks such as simulated bifurcation machines [43, 44, 45] and digital annealing [46, 47].
@@ Acknowledgements.
.1 Is diabatic enhancement a finite-size effect?
To provide additional context for our main results in sections V.1 and V.2, here we briefly discuss an observation that was already made by Shor et al. [15], namely that shorter evolution times $T$, in fact, seem to increase the success probability $P$, which they referred to as “The hare beats the tortoise”. The obvious question to ask is whether this is simply a finite-size effect at small $N$.
To conclude this Appendix, we turn to Fig. 11, which shows the ensemble statistics of the most likely final states $|\lambda\rangle$, computed for $T=128$ with the full SK ensembles as quoted in Tab. 1. For one thing, the fraction of instances for which the ground state is most likely turns out to be very small (which is, of course, to be expected by construction). More strikingly, there is a clear trend with $N$ towards higher excited states. We believe this should be interpreted as a clear indicator that diabatic enhancement becomes less effective for large systems.